OpenStudy (anonymous):

Which number in the monomial 125x18y3z25 needs to be changed to make it a perfect cube?

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Look at the components of the term to see which ones are already perfect cubes.

OpenStudy (alexandervonhumboldt2):

x shoul be in cube

Directrix (directrix):

125 = 5 ^3 x^18= ( x^6)^3 y^3 = ? z^25 = ?

Directrix (directrix):

@iam.dhue

OpenStudy (anonymous):

3 Is

Directrix (directrix):

125 = 5 ^3 x^18= ( x^6)^3 y^3 = (y)^3 z^25 = ?

Directrix (directrix):

Now, what about the power of x.

Directrix (directrix):

Divide 3 into 25. What do you get?

OpenStudy (anonymous):

.12

Directrix (directrix):

I think you did 3 divided by 25.

Directrix (directrix):

Do 25 divided by 3.